# Thermodynamics of quantum processes: An operational framework for free energy and reversible athermality

### v2

@@ VII Discussion
We develop an asymptotically reversible resource theory of athermality for quantum channels, using Gibbs-preserving superchannels as free operations. Table 1 summarizes a comparison of basic thermodynamic concepts for states (as known in the literature) and channels (all except entropy are introduced in this work). The (resource-theoretic and thermodynamic) free energy of a quantum channel quantitatively and operationally capture its distinguishability from the absolutely thermal channel with respect to a bath, thereby providing general and concrete form to the suggestions put forward in [19, 25, 31]. The discussions in [30, Section IV. A.] and [25, Sections V & VI] along with the findings in this work, provide a generic framework to extend the notion of the free energy of a quantum channel to the notion of the free energy of higher order quantum processes and the conditional free energy of a bipartite quantum processes. We will be exploring these directions rigorously and formally in future works.
@@ Appendix A Detailed Proofs
For sanwiched Rényi relative entropy $D_{\alpha}$, $\alpha=1$ is used to denote (Umegaki) relative entropy as $\lim_{\alpha\to 1}D_{\alpha}(\cdot\|\cdot)=D(\cdot\|\cdot)$. We now prove results claimed in the main content.
@@ A.1 Proof of Theorem 1
$K\geq\max\{D_{\infty}[\mathcal{N}\|\mathcal{T}^{\beta}],D_{\infty}[\mathcal{M}\|\mathcal{T}^{\beta}]\}.$ (91)
@@
The condition $\frac{1}{2}\|\mathcal{N}-\mathcal{M}\|_{\diamond}\leq\varepsilon$ implies that $\frac{1}{2}\|\mathcal{N}(\psi_{RA^{\prime}})-\mathcal{M}(\psi_{RA^{\prime}})\|_{1}\leq\varepsilon$ for all $\psi_{RA^{\prime}}$. The continuity of relative entropy in the first argument has been tightened in [90, 91]. Here we use a slightly simpler bound given in [90, Theorem 1], from which it follows that
@@
where the state $\psi^{0}_{RA^{\prime}}$ maximizes $D(\mathcal{N}(\psi_{RA^{\prime}})\|\mathcal{T}^{\beta}(\psi_{RA^{\prime}}))$ and $K_{\mathcal{N}}\geq D_{\infty}(\mathcal{N}(\psi^{0}_{RA^{\prime}})\|\mathcal{T}^{\beta}(\psi^{0}_{RA^{\prime}}))=D_{\infty}[\mathcal{N}\|\mathcal{T}^{\beta}]$. Following the similar steps for the channel $\mathcal{M}$,
@@
where $K_{\mathcal{M}}\geq D_{\infty}[\mathcal{M}\|\mathcal{T}^{\beta}]$. From Eqs. (A.1) and (96), we have
@@
To prove the inequality $\leq$”, we use the following result from [67, Lemma 38 and Proposition 41], for all $\alpha\in[1,\infty)$, for an arbitrary state $\rho_{RA^{\prime}}$, a channel $\mathcal{N}_{A^{\prime}\to A}$ , and a replacer channel $\mathcal{R}^{\tau}_{A^{\prime}\to A}$ which replaces all its input states with a state $\tau$,
@@ A.2 Proof of Lemma 1
For positive semidefinite operators $\rho\neq 0$, $\sigma$ and $\sigma^{\prime}$, if $\sigma^{\prime}\geq\sigma$, we have $D_{\alpha}(\rho\|\sigma^{\prime})\leq D_{\alpha}(\rho\|\sigma)$ for $\alpha\in[\frac{1}{2},\infty)$ ([94, Proposition 4]). Therefore, under the action of a $\mathcal{T}^{\beta}$-subpreserving superchannel, we have
@@
The invariance of $F_{\alpha}^{\beta}[\mathcal{N}]$ for $\alpha\in[\frac{1}{2},\infty)$ under Gibbs preserving unitary superchannel follows directly from [30, Proposition 1]. ∎
@@ A.4 Proof of Theorem 2
$\mathrm{Dist}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ $=\frac{1}{2}D_{H}^{\varepsilon}[\mathcal{N}\|\mathcal{T}^{\beta}],$ (137)
@@
$F^{\beta,\varepsilon}_{H}[\mathcal{N}]=\frac{2}{\beta}\mathrm{Dist}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ and $F^{\beta,\varepsilon}_{\infty}[\mathcal{N}]=\frac{2}{\beta}\mathrm{Cost}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$.
@@
The Gibbs preserving constraint on the superchannel, i.e. $\Theta_{\psi}^{\Lambda}(\mathcal{T}^{\beta})=\mathcal{R}^{\pi}$, can be written as $\operatorname{tr}(\mathcal{T}^{\beta}(\psi)\Lambda)=\frac{1}{m^{2}}$, and the conversion distance $d_{\Theta_{\psi}^{\Lambda}}((\mathcal{N},\mathcal{T}^{\beta})\rightarrow(\operatorname{id},\mathcal{R}^{\pi}))$ can be given by $1-\operatorname{tr}(\mathcal{N}(\psi)\Lambda)$. Recalling the definition of the $\varepsilon$-single-shot distillation, we have the following:
@@
$=\sup_{m,\Lambda,\psi}\{\ln m:\penalty 10000\ 1-\operatorname{tr}(\mathcal{N}(\psi)\Lambda)\leq\varepsilon,\penalty 10000\ \operatorname{tr}(\mathcal{T}^{\beta}(\psi)\Lambda)=\frac{1}{m^{2}}\}$ (140)
$=\sup_{\psi,\Lambda}\left\{-\frac{1}{2}\ln\operatorname{tr}(\mathcal{T}^{\beta}(\psi)\Lambda):\penalty 10000\ 1-\operatorname{tr}(\mathcal{N}(\psi)\Lambda)\leq\varepsilon\right\}$ (141)
$=\sup_{\psi}\left[-\inf_{\Lambda}\left\{\frac{1}{2}\ln\operatorname{tr}(\mathcal{T}^{\beta}(\psi)\Lambda):\penalty 10000\ 1-\operatorname{tr}(\mathcal{N}(\psi)\Lambda)\leq\varepsilon\right\}\right]$ (142)
$=\sup_{\psi}\frac{1}{2}D_{H}^{\varepsilon}(\mathcal{N}(\psi)\|\mathcal{T}^{\beta}(\psi))$ (143)
$=\frac{1}{2}D_{H}^{\varepsilon}[\mathcal{N}\|\mathcal{T}^{\beta}].$ (144)
Similarly, the $\varepsilon$-single-shot cost can be written as
@@
$=\inf_{m}\left\{\ln m:\penalty 10000\ \frac{1}{2}\|\mathcal{N}-\Theta(\operatorname{id})\|_{\diamond}\leq\varepsilon,\penalty 10000\ \Theta(\mathcal{R}^{\pi})=\mathcal{T}^{\beta}\right\}$ (145)
$=\inf_{m}\{\ln m:\penalty 10000\ \frac{1}{2}\|\mathcal{N}-\Theta(\operatorname{id})\|_{\diamond}\leq\varepsilon,$
$\hskip 56.9055pt\mathcal{T}^{\beta}=\frac{1}{m}\Theta(\operatorname{id})+\frac{1}{m}\sum_{i}\Theta(\mathcal{W}_{i})\}$ (146)
$=\inf_{\Theta}\left\{\ln m:\penalty 10000\ \frac{1}{2}\|\mathcal{N}-\Theta(\operatorname{id})\|_{\diamond}\leq\varepsilon,\penalty 10000\ m^{2}\mathcal{T}^{\beta}\geq\Theta(\operatorname{id})\right\}.$ (147)
Replacing $\mathcal{E}=\Theta(\operatorname{id})$ in the above relation, we have
$\mathrm{Cost}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$
$=\inf_{m}\left\{\ln m:\penalty 10000\ \frac{1}{2}\|\mathcal{N}-\mathcal{E}\|_{\diamond}\leq\varepsilon,\penalty 10000\ m^{2}\mathcal{T}^{\beta}\geq\mathcal{E}\right\}$ (148)
$=\inf_{m}\left\{\ln m:\penalty 10000\ \frac{1}{2}\|\mathcal{N}-\mathcal{E}\|_{\diamond}\leq\varepsilon,\penalty 10000\ 2\ln m\geq D_{\max}[\mathcal{E}\|\mathcal{T}^{\beta}]\right\}$ (149)
$=\inf_{m}\left\{\frac{1}{2}D_{\max}[\mathcal{E}\|\mathcal{T}^{\beta}]:\penalty 10000\ \|\mathcal{N}-\mathcal{E}\|_{\diamond}\leq\varepsilon\right\}$ (150)
@@ A.5 SDP for the max-free energy
It follows from the fact that the max-relative entropy between channels is an SDP where strong duality holds [49], i.e., optimal solution of the primal and dual problems are the same.
@@ A.7 Proof of Proposition 3
$\mathrm{Cost}^{\sqrt{\varepsilon}}(\mathcal{N},\mathcal{T}^{\beta})$ $\leq\mathrm{Dist}^{1-\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})+\frac{1}{2}\ln\left(\frac{1}{1-\varepsilon}\right).$ (156)
@@
Note that given a channel $\mathcal{M}_{A^{\prime}\to A}$ s.t. $\mathcal{M}\in\mathcal{B}^{\epsilon}[\mathcal{N}]$ implies that $\mathcal{M}(\psi)\in\mathcal{B}^{\varepsilon}(\mathcal{N}(\psi))$ for all $\psi\in\operatorname{St}(RA^{\prime})$. Therefore, the smoothed max-relative channel entropy between any channel $\mathcal{N}$ and $\mathcal{P}$ satisfies the following inequality [cf. 38, Lemma 1],
@@
The proof follows since $\mathrm{Cost}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})=\frac{1}{2}D_{\infty}^{\varepsilon}[\mathcal{N}\|\mathcal{T}^{\beta}]$ and $\mathrm{Dist}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})=\frac{1}{2}D_{\rm H}^{\varepsilon}[\mathcal{N}\|\mathcal{T}^{\beta}]$. ∎

### v3

@@ VII Discussion
We develop an asymptotically reversible resource theory of athermality for quantum channels, using Gibbs-preserving superchannels as free operations. Table 1 summarizes a comparison of basic thermodynamic concepts for states (as known in the literature) and channels (all except entropy are introduced in this work). The (resource-theoretic and thermodynamic) free energy of a quantum channel quantitatively and operationally capture its distinguishability from the absolutely thermal channel with respect to a bath, thereby providing general and concrete form to the suggestions put forward in [71, 85, 25]. The discussions in [23, Section IV. A.] and [85, Sections V & VI] along with the findings in this work, provide a generic framework to extend the notion of the free energy of a quantum channel to the notion of the free energy of higher order quantum processes and the conditional free energy of a bipartite quantum processes. We will be exploring these directions rigorously and formally in future works.
@@ Appendix A Detailed Proofs
For sanwiched Rényi relative entropy $D_{\alpha}$, $\alpha=1$ is used to denote (Umegaki) relative entropy as $\lim_{\alpha\to 1}D_{\alpha}(\cdot\|\cdot)=D(\cdot\|\cdot)$. Additionally, given two completely positive maps $\mathcal{N}$ and $\mathcal{M}$, we say $\mathcal{N}\geq\mathcal{M}$ if the linear map $\mathcal{N}$-$\mathcal{M}$ is also completely positive. We now prove the results claimed in the main content.
@@ A.1 Proof of Theorem 1
$K\geq\max\{D_{\infty}(\mathcal{N}(\psi^{0}_{RA^{\prime}})\|\mathcal{T}^{\beta}(\psi^{0}_{RA^{\prime}})),D_{\infty}(\mathcal{M}(\phi^{0}_{RA^{\prime}})\|\mathcal{T}^{\beta}(\phi^{0}_{RA^{\prime}}))\}.$ (92)
where $\psi^{0}_{RA^{\prime}}$ maximizes $D(\mathcal{N}(\psi_{RA^{\prime}})\|\mathcal{T}^{\beta}(\psi_{RA^{\prime}}))$ and $\phi^{0}_{RA^{\prime}}$ maximizes $D(\mathcal{M}(\psi_{RA^{\prime}})\|\mathcal{T}^{\beta}(\psi_{RA^{\prime}}))$ respectively.
@@
The condition $\frac{1}{2}\|\mathcal{N}-\mathcal{M}\|_{\diamond}\leq\varepsilon$ implies that $\frac{1}{2}\|\mathcal{N}(\psi_{RA^{\prime}})-\mathcal{M}(\psi_{RA^{\prime}})\|_{1}\leq\varepsilon$ for all $\psi_{RA^{\prime}}$. The continuity of relative entropy in the first argument has been tightened in [12, 3]. Here we use a slightly simpler bound given in [12, Theorem 1], from which it follows that
@@
where the state $\psi^{0}_{RA^{\prime}}$ maximizes $D(\mathcal{N}(\psi_{RA^{\prime}})\|\mathcal{T}^{\beta}(\psi_{RA^{\prime}}))$ and $K_{\mathcal{N}}\geq D_{\infty}(\mathcal{N}(\psi^{0}_{RA^{\prime}})\|\mathcal{T}^{\beta}(\psi^{0}_{RA^{\prime}}))$. Following the similar steps for the channel $\mathcal{M}$,
@@
where $K_{\mathcal{M}}\geq D_{\infty}(\mathcal{M}(\phi^{0}_{RA^{\prime}})\|\mathcal{T}^{\beta}(\phi^{0}_{RA^{\prime}}))$ where the state $\phi^{0}_{RA^{\prime}}$ maximizes $D(\mathcal{M}(\psi_{RA^{\prime}})\|\mathcal{T}^{\beta}(\psi_{RA^{\prime}}))$. From Eqs. (A.1) and (97), we have
@@
To prove the inequality $\leq$”, we use the following result from [92, Lemma 38 and Proposition 41], for all $\alpha\in[1,\infty)$, for an arbitrary state $\rho_{RA^{\prime}}$, a channel $\mathcal{N}_{A^{\prime}\to A}$ , and a replacer channel $\mathcal{R}^{\tau}_{A^{\prime}\to A}$ which replaces all its input states with a state $\tau$,
@@ A.2 Proof of Lemma 1
For positive semidefinite operators $\rho\neq 0$, $\sigma$ and $\sigma^{\prime}$, if $\sigma^{\prime}\geq\sigma$, we have $D_{\alpha}(\rho\|\sigma^{\prime})\leq D_{\alpha}(\rho\|\sigma)$ for $\alpha\in[\frac{1}{2},\infty)$ ([70, Proposition 4]). Therefore, under the action of a $\mathcal{T}^{\beta}$-subpreserving superchannel, we have
@@
The invariance of $F_{\alpha}^{\beta}[\mathcal{N}]$ for $\alpha\in[\frac{1}{2},\infty)$ under Gibbs preserving unitary superchannel follows directly from [23, Proposition 1]. ∎
@@ A.4 Proof of Theorem 2
$\mathrm{Dist}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ $=\frac{1}{2}D_{H}^{\varepsilon^{2}}[\mathcal{N}\|\mathcal{T}^{\beta}],$ (138)
@@
$F^{\beta,\varepsilon}_{H}[\mathcal{N}]=\frac{2}{\beta}\mathrm{Dist}^{\sqrt{\varepsilon}}(\mathcal{N},\mathcal{T}^{\beta})$ and $F^{\beta,\varepsilon}_{\infty}[\mathcal{N}]=\frac{2}{\beta}\mathrm{Cost}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$.
@@
The Gibbs preserving constraint on the superchannel, i.e. $\Theta_{\psi}^{\Lambda}(\mathcal{T}^{\beta})=\mathcal{R}^{\pi}$, can be written as $\operatorname{tr}(\mathcal{T}^{\beta}(\psi)\Lambda)=\frac{1}{m^{2}}$, and the conversion distance $d_{\Theta_{\psi}^{\Lambda}}((\mathcal{N},\mathcal{T}^{\beta})\rightarrow(\operatorname{id},\mathcal{R}^{\pi}))$ is given by $\sqrt{1-\operatorname{tr}(\mathcal{N}(\psi)\Lambda)}$. Recalling the definition of the $\varepsilon$-single-shot distillation, we have the following:
@@
$=\sup_{m,\Lambda,\psi}\{\ln m:\penalty 10000\ \sqrt{1-\operatorname{tr}(\mathcal{N}(\psi)\Lambda)}\leq\varepsilon,\penalty 10000\ \operatorname{tr}(\mathcal{T}^{\beta}(\psi)\Lambda)=\frac{1}{m^{2}}\}$ (141)
$=\sup_{\psi,\Lambda}\left\{-\frac{1}{2}\ln\operatorname{tr}(\mathcal{T}^{\beta}(\psi)\Lambda):\penalty 10000\ 1-\operatorname{tr}(\mathcal{N}(\psi)\Lambda)\leq\varepsilon^{2}\right\}$ (142)
$=\sup_{\psi}\left[-\inf_{\Lambda}\left\{\frac{1}{2}\ln\operatorname{tr}(\mathcal{T}^{\beta}(\psi)\Lambda):\penalty 10000\ 1-\operatorname{tr}(\mathcal{N}(\psi)\Lambda)\leq\varepsilon^{2}\right\}\right]$ (143)
$=\sup_{\psi}\frac{1}{2}D_{H}^{\varepsilon^{2}}(\mathcal{N}(\psi)\|\mathcal{T}^{\beta}(\psi))$ (144)
$=\frac{1}{2}D_{H}^{\varepsilon^{2}}[\mathcal{N}\|\mathcal{T}^{\beta}].$ (145)
Now, to derive the $\varepsilon$-single-shot cost, we first calculate the conversion distance:
$d_{\mathrm{GP}}((\operatorname{id}_{m}\mathcal{R}^{\pi})\rightarrow(\mathcal{N},\mathcal{T}^{\beta}))$
$\qquad=\min_{\Theta}\left\{P[\mathcal{N},\Theta(\operatorname{id}_{m})]\penalty 10000\ :\penalty 10000\ \Theta(\mathcal{R}^{\pi})=\mathcal{T}^{\beta}\right\}$ (146)
$\qquad=\min_{\Theta}\left\{P[\mathcal{N},\Theta(\operatorname{id}_{m})]\penalty 10000\ :\right$.
$\hskip 71.13188pt\left.\mathcal{T}^{\beta}=\frac{1}{m^{2}}\Theta(\operatorname{id}_{m})+\frac{1}{m^{2}}\Theta(\operatorname{id}_{m}^{\perp})\right\}$ (147)
$\qquad=\min_{\mathcal{E}}\left\{P[\mathcal{N},\mathcal{E}]\penalty 10000\ :\penalty 10000\ m^{2}\mathcal{T}^{\beta}\geq\mathcal{E}\right\}$ (148)
$\qquad=\min_{\mathcal{E}}\left\{P[\mathcal{N},\mathcal{E}]\penalty 10000\ :\penalty 10000\ 2\ln m\geq D_{\infty}[\mathcal{E}\|\mathcal{T}^{\beta}]\right\}.$ (149)
where we have replaced $\mathcal{E}=\Theta(\operatorname{id}_{m})$. Note that $\mathcal{E}$ can be any channel such that $m^{2}\mathcal{T}^{\beta}\geq\mathcal{E}$, and is not constrained by the Gibbs preserving property of $\Theta$. This is because the quantity that we are optimizing is a function of the variable $\Theta(\operatorname{id}_{m})$, and the range of this variable is fully captured by the constraint $m^{2}\mathcal{T}^{\beta}\geq\Theta(\operatorname{id}_{m})$. The $\varepsilon$-single-shot cost can be written as
@@
$=\inf_{m}\left\{\ln m:\penalty 10000\ d_{\mathrm{GP}}((\operatorname{id}_{m}\mathcal{R}^{\pi})\rightarrow(\mathcal{N},\mathcal{T}^{\beta}))\leq\varepsilon\right\}$ (150)
$=\inf_{m}\{\ln m:\penalty 10000\ P[\mathcal{N},\mathcal{E}]\leq\varepsilon,\penalty 10000\ 2\ln m\geq D_{\infty}[\mathcal{E}\|\mathcal{T}^{\beta}]\}$ (151)
$=\inf_{\mathcal{E}}\left\{\frac{1}{2}D_{\max}[\mathcal{E}\|\mathcal{T}^{\beta}]:\penalty 10000\ P[\mathcal{N},\mathcal{E}]\leq\varepsilon\right\}$ (152)
@@ A.5 SDP for the max-free energy
It follows from the fact that the max-relative entropy between channels is an SDP where strong duality holds [58], i.e., optimal value of the primal and dual problems are the same.
@@ A.7 Proof of Proposition 3
$\mathrm{Cost}^{\sqrt{\varepsilon}}(\mathcal{N},\mathcal{T}^{\beta})$ $\leq\mathrm{Dist}^{\sqrt{1-\varepsilon}}(\mathcal{N},\mathcal{T}^{\beta})+\frac{1}{2}\ln\left(\frac{1}{1-\varepsilon}\right).$ (158)
@@
Note that given a channel $\mathcal{M}_{A^{\prime}\to A}$ s.t. $\mathcal{M}\in\mathcal{B}^{\epsilon}[\mathcal{N}]$ implies that $\mathcal{M}(\psi)\in\mathcal{B}^{\varepsilon}(\mathcal{N}(\psi))$ for all $\psi\in\operatorname{St}(RA^{\prime})$. Therefore, the smoothed max-relative channel entropy between any channel $\mathcal{N}$ and $\mathcal{P}$ satisfies the following inequality [cf. 4, Lemma 1],
@@
The proof follows since $\mathrm{Cost}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})=\frac{1}{2}D_{\infty}^{\varepsilon}[\mathcal{N}\|\mathcal{T}^{\beta}]$ and $\mathrm{Dist}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})=\frac{1}{2}D_{\rm H}^{\varepsilon^{2}}[\mathcal{N}\|\mathcal{T}^{\beta}]$. ∎